OpenStudy (anonymous):

In a triangle ABC, angle A is congruent to angle C, AB= 10x-7, BC=2x+33, and AC= 4x-6. Find x.

OpenStudy (anonymous):

(10x - 7) + (2x + 33) + (4x - 6) = 180 -- the angles in an isolesces triangle = 180 degrees combine like terms 16x + 20 = 180 16x = 180 - 20 16x = 160 x = 160/16 x = 10 I am not the best at geometry, so you might better get a second opinion

OpenStudy (anonymous):

typo * isosceles

OpenStudy (anonymous):

am I way off on this ?

OpenStudy (anonymous):

AB = 10x - 7 : 10(10) - 7 : 100 - 7 = 93 BC = 2x + 33 : 2(10) + 33 : 20 + 33 = 53 AC = 4x - 6 : 4(10) - 6 : 40 - 6 = 34 93 + 53 + 34 = 180

OpenStudy (anonymous):

but I am not getting two of the same angles...

OpenStudy (ybarrap):

 $$ \cfrac{\sin C}{10x-7}=\cfrac{\sin A}{2x+33}\\ $$ But, \(\sin C=\sin A\), so by the Law of Sines: $$ 1=\cfrac{10x-7}{2x+33}\\ 2x+33=10x-7\\ 2x-10x=-7-33\\ -7x=-40\\ x=\cfrac{40}{7} $$

OpenStudy (anonymous):

I thought A and C were the base angles making it an isosceles triangle ? I am totally off track here.....

OpenStudy (ybarrap):

They are -- A and C are the base angles making an isosceles triangle. So the segments opposite them are equal: \(2x+33=4x-6\).

OpenStudy (anonymous):

so if 40/7 was plugged in for x, all 3 angles would equal 180 ??

OpenStudy (ybarrap):

AB, AC and BC are all lengths from each of the vertices, not angles. If they had been angles, the problem would have stated angle A, angle B and angle C when referencing them.